OpenStudy (anonymous):

The diameter of the circle above is 38 in. What is the circumference of the circle? (Use = 3.14.)

OpenStudy (jdoe0001):

\(\textit{circumference of a circle}=2\pi r\qquad \begin{cases} r=radius=\frac{diameter}{2} \end{cases}\)

OpenStudy (anonymous):

So whats the answer?

OpenStudy (jdoe0001):

if the diameter is 38, then the radius is half that use that :)

OpenStudy (anonymous):

These are my options 2,267.08 in 1,133.54 in 59.66 in 119.32 in

OpenStudy (jdoe0001):

so... what would you get for \(\bf 2\pi r?\)

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

Can you help me on this one instead The diameter of the circle above is 6 cm. What is the circumference of the circle? (Use = 3.14.) 9.42 cm 18.84 cm 28.26 cm 56.52 cm

OpenStudy (jdoe0001):

\(\bf 2\pi r\implies 2\cdot \pi \cdot r\) is the same as the one before just get the radius, and use it in in \(\bf 2\pi r\implies 2\cdot \pi \cdot r\)

OpenStudy (jdoe0001):

notice, you're asked to use \(\bf (\pi = 3.14)\)

OpenStudy (anonymous):

Thanks!!!!!! I got iit right how bout this one........ A circle has an area of 200.96 units2 and a circumference of 50.24 units. If the radius is 8 units, what can be said about the relationship between the area and the circumference? (Use 3.14 for .) The ratio of the area to the circumference is equal to the square root of the radius. The ratio of the area to the circumference is equal to the radius squared. The ratio of the area to the circumference is equal to twice the radius. The ratio of the area to the circumference is equal to half the radius.

OpenStudy (anonymous):

Nvm mind about that how bout this one instead..For the given circle, which of the following is true regarding its circumference and area? The circumference is larger, and it is two times the area. The circumference is larger, and it is four times the area. The area is larger, and it is three times the circumference. The area is larger, and it is four times the circumference.

OpenStudy (jdoe0001):

you may want to post anew, more eyes

OpenStudy (anonymous):

Do you know the formula To find the area of a circle